# Jacobian-Enhanced Neural Networks

### v2

@@ V Example Application: Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural networks, to optimize some slow-running function of interest, rather than directly using the function itself. This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[26]. The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization. This will be accomplished by solving the following optimization problem:
@@
where is a surrogate-model approximation to the Rosenbrock [27] function, which is used as an optimization benchmark because it exhibits a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization:
@@
Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3b, where optimization histories are shown for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished). Training data is not shown to avoid clutter, but was generated by sampling the true response on a regular grid plus 100 latin hypercube points. The neural network training histories are also included in Fig. 3a, in order to verify that both neural networks converged to the same degree.
V-A Results Without Polishing
@@
Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed: neither JENN nor NN reach the true location of this optimum situated at . Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer. These effects are less pronounced for JENN, but the optimizer gets trapped nonetheless.
In hindsight, this finding is not surprising because neural networks learn by minimizing prediction error (recall Eq. III-D). Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting flat regions in the vicinity of a zero optimum, where prediction error penalties are negligible. This situation is at odds with the needs of gradient-based optimization which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
@@
For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate. However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design where it is imperative that the final airfoil is free of undulations and non-smooth contours. An imperative research question therefore arises: how can these competing needs be reconciled?
V-B Results With Polishing
@@
This situation can be ameliorated through “polishing.” Given a trained model, this approach consists of train it further after magnifying regions of interest. This is accomplished by judiciously setting to prioritize specific training points belonging to that subspace. For example, flat regions could be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on such as:
@@
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are whereas near those points . This has the desired effect of disproportionately emphasizing regions of small slope during training by up to a factor of 1000 at the expense of de-emphasizing other regions, which is why it is necessary to start from an already trained model.
Final results are shown in Fig. 3b for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly. Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
VI Noisy Partials
@@
We now consider the effect of noisy partials. As mentioned in the introduction, many practical design applications rely on legacy software suites, such as missile DATCOM [28] during early design phases, for which analytical or automatically differentiated partials are not available. In this case, the only recourse for computing derivatives is finite difference. However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts [29]. It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference is unavoidable?
@@
The answer to this question is shown by example in Fig. 4 using the Rastrigin function as a testbed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural network (gradient-enhancement off). Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
@@
Finally, these results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials. For most applications, the hyperparameter should therefore be treated as a binary scalar to turn gradient-enhancement on () or off (); except in special situations which call for distribution (rather than a scalar) such as polishing, as previously illustrated, or incomplete partials. The latter is likely to be a common scenario where partials are not available for all inputs, in which case can be used to turn gradient-enhancement on/off depending on whether a partial is missing or not (i.e. resulting in a tensor of zeros and ones).
VII Conclusion
In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions. It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place. It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization. However, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero. It was further shown that this shortcoming can be overcome through polishing, which can be used to prioritize salient regions of interest through judicious use of hyperparameters. This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it. The ideas and findings of this paper have been published as an open-source library to allow readers to validate the results, and hopefully assist them in their surrogate modeling efforts.
Acknowledgement
@@
The supporting code for this work used the exercises by Professor Andrew Ng [23] as a starting point. It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied). The author would therefore like to thank Professor Andrew Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy to understand form.
VIII Key Equations in Vectorized Form
@@ VIII-A Data Structures
VIII-B Forward Propagation (Eq. 3)
VIII-C Parameter Update (Eq. 8)
VIII-D Backward Propagation (Eqs. 15–16, 17, 20)

### v3

@@ 4.3 Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural network, to optimize some slow-running function of interest, rather than directly using the function itself. This use case typically arises when the response of interest takes too long to evaluate relative to the time available for analysis. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory (Sanchez2018). The purpose of this section is to quantify the benefit of JENN relative to regular neural networks in the context of gradient-based optimization. This will be accomplished by solving the following, constrained optimization problem:
@@
where is a surrogate-model approximation to the Rosenbrock function (Rosenbrock1960), which is often used as a benchmark in optimization because the optimum lies in a “shallow valley” with near-zero slope that stalls convergence:
@@
Results are shown in Fig.4, which compares optimization convergence histories for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished). Training data is not shown to avoid clutter, but was generated by sampling the true response on a regular grid plus 100 Latin hypercube points.
4.3.1 Results Without Polishing
@@
Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not the case. Neither JENN nor NN reach the true location of this optimum situated at with a value of zero. Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, likely within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer. This behavior is the Achille’s heel of surrogate-based optimization methods. These effects are less pronounced for JENN, but the optimizer gets trapped nonetheless. This finding is not surprising whenever the optimum lies in a flat region and approaches zero, since magnitudes are so small they cannot be distinguished from prediction error.
@@
For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate. However, there are situations where optimization accuracy is desirable if not critical, such as aerodynamic shape optimization. In general, surrogate-based optimization tends to struggle any time the magnitude of the slopes in the region of interest lie within prediction noise. This situation is at odds with the needs of gradient-based optimization which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement. An imperative research question therefore arises: how can these competing needs be reconciled?
4.3.2 Results With Polishing
@@
This situation can be ameliorated through “polishing” which is uniquely afforded by this framework. Given a trained model, this approach consists of training it further after magnifying regions of interest. This is accomplished by judiciously setting to prioritize specific training points belonging to that subspace. Specifically, flat regions can be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on for example:
@@
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are whereas near those points . This has the desired effect of disproportionately emphasizing regions of small slope during training up to a factor of 1000 at the expense of deprioritizing other regions, which is why it is necessary to start from an already trained model. Results are shown in Fig. 4 for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly.
4.4 Noisy Partials
@@
We now consider the effect of noisy partials. Many practical design applications rely on legacy software suites during early design phases, such as missile DATCOM (DATCOM), for which analytical or automatically differentiated partials are not available. In this case, the only recourse for computing derivatives is finite difference. However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts (MartinsHwang2013). It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference might be unavoidable?
@@
The answer to this question is shown by example in Fig. 5 using the Rastrigin function as a testbed once again, where corresponds to baseline results for a standard neural network (gradient-enhancement turned off). Partial derivatives of the training data were obtained using finite difference with progressively increasing step sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
@@
These results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials. For most applications, the hyperparameter should therefore be treated as a binary scalar variable to turn gradient-enhancement on () or off (); except in special situations such as incomplete partials or polishing, as previously illustrated, which call for a distribution of in that case. Incomplete partials refers to the scenario where partials are only available for some inputs, in which case can be used to turn gradient-enhancement off for the missing partials, i.e. resulting in a tensor of zeros and ones as illustrated next in airfoil application section.
5 Application: Airfoil Shape Optimization
This section quantifies the benefit of JENN for aerodynamic shape optimization through applied example. This is accomplished by optimizing the shape of a NACA 0012 airfoil in inviscid, transonic flow and comparing results to status quo neural networks.
5.0.1 Problem Description
To ensure reproducibility, the open-source software SU2 (Economon2016) was selected for computational fluid dynamics. The design problem is taken directly from their website555https://su2code.github.io/tutorials/Inviscid_2D_Unconstrained_NACA0012/, modified to hold lift coefficient fixed (instead of angle of attack), and airfoil geometry is parameterized using fourteen Hicks-Henne “bump” functions (Hicks1978). The goal is to minimize drag at some desired flight condition, specified by Mach number and lift coefficient :
| Minimize: | (27) | |||
| With Respect To: | (28) | |||
| Subject To: | (29) |
where denote the Hicks-Henne shape parameters. The flight condition parameters determine the function , which effectively defines a different instance of the optimization problem. Each instance took several minutes to solve using adjoint methods on a modern laptop.
5.0.2 Statistical Regression
Our goal is to accelerate numerical optimization by developing an optimal proxy, which can be used to near instantaneously shape an airfoil at any flight condition. In the proposed approach, this is accomplished by replacing the objective function by an equivalent approximation , taken to be a neural network. To do so, one must therefore sample a 16-dimensional input space composed of 14 shape variables and two flight conditions and to generate training data. Sample results are shown in Fig. 6 for a subset of the design space. A total of 10,000 Latin hypercube samples were evaluated.
Upon collecting samples, the data was used to determine how many training points were needed to achieve good prediction accuracy using JENN and regular Neural Networks (NN). This was accomplished by holding back a subset of the data for training and using the rest for testing. The neural net architecture (5 hidden layers of 16 neurons each) and hyperparameter settings were identical for both neural networks, except for gradient enhancement which was turned on or off accordingly. The results are shown in Fig. 7 using test data to plot the R-squared value as function of the number of training samples. Results unambiguously show that JENN achieves the same prediction accuracy of 99% using five times fewer samples, re-affirming the conclusion by other authors that gradient-enhanced methods reduce the amount of data required for good prediction, even on practical engineering problems.
Using the results of Fig. 7, it is possible to identify the minimum number of points needed to train an accurate neural net under both paradigms for this problem, as indicated by the arrows. Associated convergence properties and goodness of fit were verified to be satisfactory but not reported for conciseness, except for Table 1 which compares the R-squared value of the partials in order to emphasize that JENN predicts them more accurately.
| NN | -0.69 | 0.79 | 0.79 | 0.68 | 0.32 | 0.26 | 0.32 | -0.95 | 0.73 | 0.80 | 0.77 | 0.59 | 0.49 | -0.10 |
| JENN | 0.37 | 0.90 | 0.89 | 0.78 | 0.59 | 0.18 | 0.34 | 0.29 | 0.85 | 0.88 | 0.87 | 0.77 | 0.59 | 0.13 |
Corresponding sensitivity profiles are shown in Fig. 8 for all variables, in order to show the topological differences that resulted from gradient enhancement. Overall, it was found that JENN and NN exhibit excellent agreement in the subsonic regime, but that discrepancies start to appear in transonic flow at higher Mach numbers, where shock waves are present, presumably creating stronger sensitivity to shape variables that control the region where shocks tend to appear. As a result, the surrogate model learns to predict sensitivities more accurately in some regions than others, which can be seen from the prediction profiles. Although there is good agreement in general, notice that NN slopes occasionally point in a different direction in regions that tend to be flatter. This could have subtle consequences during optimization.
Finally, it should be noted that SU2 did not provide partial derivatives for or , which implies incomplete gradient information. This is particularly noteworthy because most engineering design applications will fall in this scenario: partials are only available for some variables. Yet, by virtue of the framework afford by JENN, such missing information is easily overcome by simply setting for the missing partials. This represents a simple but important practical improvement over similar reported work.
5.0.3 Surrogate-Based Optimization
The surrogate models developed in the previous subsection were substituted for in the optimization problem described by Eqs. 27–29, which was solved using the SLSQP algorithm from Scipy (2020SciPy-NMeth). Fig. 4 compares final solutions against the initial geometry using different approaches. All solutions successfully recovered a supercritical airfoil shape, as expected, but adjoint-based optimization took several minutes, whereas surrogate-based optimization took less than a second, setting the cost of generating training data aside. However, there are some observable differences. In particular, notice the presence of a weak but unwanted secondary shock wave in the answer obtained using a regular NN. This is necessarily attributed to small differences in predicted partial derivatives, since that is the only difference between JENN and a regular NN. This underlines the foreseeable conclusion that gradient-enhanced methods ameliorate surrogate-based optimization somewhat whenever gradient-descent algorithms are used. Overall, predictions remained comparable, but the cost of generating data was significantly less for JENN.
6 Conclusion
In summary, this research developed JENN, which are densely connected multi-layer perceptrons that predict gradients more accurately, successfully validated the implementation against canonical test functions, and applied it to a practical airfoil shape optimization problem. All results support the general consensus that gradient-enhanced methods require less training data than their status quo neural network counterparts and achieve better accuracy, provided the cost of obtaining training data partials is not prohibitive in the first place.
It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization, but that surrogate-based methods are more prone to getting trapped in artificial local minima, whenever the true optimum lies in flat regions and approaches zero. This was apparent for the Rosenbrock test function, but also for the airfoil design problem where the regular NN solution exhibited a weak secondary shockwave that was not present with JENN. Overall, JENN tends to achieve optimal solutions that are in better agreement with the true optimum.
Furthermore, through a process dubbed polishing, uniquely afforded by this framework, JENN enables salient regions of interest to be magnified through judicious use of hyperparameters. This feature provides an additional knob for improving surrogate-based optimization accuracy in situations that call for it, as demonstrated with the Rosenbrock example. The same mechanism also enables missing partials to be ignored during training by setting for those partials, such that gradient-enhancement can be leveraged even in situations when information is incomplete. This is likely to be a common scenario for many engineering design problems, as was the case for the airfoil design example where and were missing. This simple mechanism is likely to be of important practical value.
In closing, despite its advantages over status quo neural networks, JENN does not overcome the curse of dimensionality; although it alleviates it by reducing the number of samples needed. Sampling remains necessary to generate training data, and so the practical utility of gradient-enhanced methods is likely to be restricted to problems where intrinsic dimensionality is either small to start with or can be recovered using elaborate dimensionality reduction techniques (BerguinMavris2015; BerguinRancourtMavris2015; Constantine2015). To be adopted at industrial scales, new techniques are needed to avoid high-dimensional sampling whilst retaining the benefit of optimal proxies for accelerating numerical optimization.
Acknowledgments
@@
The supporting code for this work used the exercises by Professor Andrew (Ng2017) as a starting point, but fundamentally change the formulation to handle Jacobian enhancement. The primary author, who was then a student, would therefore like to thank Professor Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy way to understand.
Appendix A Key Equations in Vectorized Form
@@ A.1 Data Structures
A.2 Forward Propagation (Eq. 3)
A.3 Parameter Update (Eq. 8)
A.4 Backward Propagation (Eqs. 15–16, 17, 20)